OpenStudy (anonymous):

Steps on how to establish this identity: (sinθ-cosθ+1)/(sinθ+cosθ-1) = (sinθ+1)/cosθ

OpenStudy (e.mccormick):

What have you tried?

OpenStudy (e.mccormick):

\[\frac{\sin\theta-\cos\theta+1}{\sin\theta+\cos\theta-1}=\frac{\sin\theta+1}{\cos\theta}\]

OpenStudy (anonymous):

I tried writing it as\[\frac{ \sinθ-(\cosθ-1) }{ \sin+ (\cosθ-1) } \times \frac{ \sinθ-(\cosθ-1) }{ \sinθ-(\cosθ-1}\] But then I realized that I was just crossing out the numerator and denominator... so now I'm stuck.

OpenStudy (e.mccormick):

Yah, it is not a friendly one where something is obvious. If you want to simplify the left, the sum to product things might work, but I am going to try something else. I think I rememer seeing something like this, but not 100% sure.

OpenStudy (fibonaccichick666):

Can you factor out a \[sin\theta -1\]?

OpenStudy (e.mccormick):

OK. the start was right. It gets there.

OpenStudy (e.mccormick):

It is ugly tho.

OpenStudy (fibonaccichick666):

I'll say

OpenStudy (dan815):

u wanna know a silly way of doing all these problems

OpenStudy (e.mccormick):

Yah, you have to multiply that all out, which eventually lets you cancel a 2 out of everything. Then there is some grouping, and some other cancelation.

OpenStudy (dan815):

if the integral or derivative of LHS = RHS then the 2 functions are equal argument

OpenStudy (dan815):

just thougght i throw that in there >_>

OpenStudy (e.mccormick):

Not exactly useful for someone in trig.

OpenStudy (dan815):

maybe make it integral + derivative of LHS = RHS to make it more concrete

OpenStudy (dan815):

lol

OpenStudy (fibonaccichick666):

Still not helpful for someone in trig

OpenStudy (fibonaccichick666):

haha

OpenStudy (dan815):

i know lol but emccormick got this covered!!

OpenStudy (e.mccormick):

OK. I am going to write out my version. What you started with will work. It is one o those things where it is a little messy, but it did get there. I think I have collapsed some steps, which is a no-no for a proof, so you may need to expand some back out and fill in the gaps.

OpenStudy (fibonaccichick666):

^go @e.mccormick you got this!!!!

OpenStudy (aravindg):

really ??

OpenStudy (aravindg):

I dont understand why you are asking this question ....Its exactly similar to the last one we worked on !

OpenStudy (anonymous):

I did it the same method! ^^ look above. I'm stuck!

OpenStudy (aravindg):

why ? first of all divide by cos theta on both sides .You get the previous question , then I have worked it for you

OpenStudy (aravindg):

rhs can be split as sec theta and tan theta

OpenStudy (e.mccormick):

Well, this start does work. It makes it a bunch of algebra.

OpenStudy (fibonaccichick666):

When confirming an Identity, you cannot work with both sides... You must choose one and get the other.

OpenStudy (dan815):

just do this

OpenStudy (aravindg):

(sinθ-cosθ+1)/(sinθ+cosθ-1) divide by cos theta on numerator and denominator

OpenStudy (aravindg):

You will see that its exactly the previous prove that question

OpenStudy (e.mccormick):

However, taking on the RHS of an ID is always worth thinking about. It can be a very good solution. As long as you stick on the right... or just use it as a way of seeing things to aim for on the left, sort of a guide, there are lots of tricks that can be done.

OpenStudy (dan815):